# Guiding polarizable particles in multi-hole Gaussian beams

### v1

@@ Abstract
The present paper discusses certain special Gaussian beams that, thanks to some polynomial prefactors, have uniquely designed holes in the irradiance. Such holes, or rather tubes, can constitute potential valleys for negatively polarizable particles, providing the possibility of guiding several objects of that kind, each along its own trajectory. The mechanism of creating these holes by interference of Gaussian beams which exhibit orbital angular momentum is discussed, and then the trajectories of particles moving in such a wave are numerically calculated. As it turns out, these particles, performing transverse oscillations, follow the designed tunnels of low irradiance. On the contrary, for particles with positive polarizability these areas are inaccessible.
@@ I Introduction
The original optical tweezers operated due to gradient forces pulling particles of positive polarizability into areas of high wave intensity. In the similar way, the irradiance “holes” can serve as traps or guide lines for objects of negative polarizability such as atoms in the blue-detuned beams [16, 17, 18]. The identical effect is owed to the ponderomotive force acting on charged particles, such as electrons, originating from the inhomogeneous circularly polarized wave [19].
@@ II Description of multi-hole Gaussian beams
In the above formulas (9) and (10) the complex coordinate $\xi=\xi_{x}+i\xi_{y}$ is introduced and should be distinguished from $|\bm{\xi}|=\sqrt{\xi_{x}^{2}+\xi_{y}^{2}}=|\xi|$. Result (10) is well known [45]. In particular the choice of $\tilde{\psi}(s)=\mathrm{const}$ leads to the fundamental GB, and $\tilde{\psi}(s)=\mathrm{const}\cdot s^{n}$ corresponds to the GB of vorticity $n$.
@@
$\left\{\begin{array}[]{l}|\bm{\xi}|=\displaystyle\frac{p_{0}}{\beta}\sqrt{1+\zeta^{2}},\\ \phi(\zeta)=\phi_{0}+\arctan\zeta.\end{array}\right.$ (11)
Since all possible zeros follow synchronized, identical paths, they never merge, their number remains constant along the beam, and each develops its own nodal line as the value of $\zeta$ increases (naturally in the opposite direction, i.e. for $\zeta<0$ as well).
@@
The role of the parameter $\beta$ will become clear later. This form indicates that the interference of exactly two Gaussian beams is dealt with: that of order 00:
@@
Beam (13) bears the vortex topological charge $n$ and when encircling the axis $\zeta$ the value of the phase increases by $2\pi n$ which means that it assumes $n$ times the same values (if reduced to the interval $[0,2\pi[$). Thus, on a circle (for $\zeta=\mathrm{const}$) of radius $\sqrt{1+\zeta^{2}}/\beta$ (in units of $w_{0}$) a completely destructive interference with $\tilde{\psi}_{0}$ occurs exactly $n$ times. The $n$th degree vortex “spreads” into $n$ individual vortices, uniformly distributed, as is obvious from the distribution of the $n$th complex roots of the unity. Consequently, there appear $n$ “holes” in the wave intensity distribution in the perpendicular plane, as shown in Fig. 1 for $n=3$. In the subsequent diagrams (a)-(h) performed for increasing values of $\zeta$, the diffraction of the wave and the twisting of the whole pattern can be seen, in agreement with (11).
@@
Figures 3 and 4 demonstrate the same effect for $n=5$. The existence of five distinct roots of unity yields five “tubes” of vanishing irradiance that can eventually be exploited. A slightly reduced value for the parameter $\beta$ has been chosen in this case in order to avoid merging areas of low energy density.
@@ III Guiding of particles
As is well known, inhomogeneities in the intensity of the wave, and hence in the electric field, can be exploited to trap and guide neutral particles, for instance atoms, which undergo polarization in external fields. This has become the basis for the operation of the so-called optical tweezers [12, 13, 14, 15], as mentioned in Introduction.
Figure 7: The trajectories of three particles of negative polarizability placed in the “holes” of the beam in question for $n=3$, viewed from the side and from above. Straight lines represent zero irradiance tubes, which undergo diffraction. The values of beam parameters are the same as in Fig. 1, and $\gamma=6\cdot 10^{-2}$ and $\tilde{\gamma}=10^{-2}$.
Let us denote with $\alpha$ the atomic polarizability, which in general depends on the driving frequency, and with $d$ the induced dipole moment. Then
@@
From the theory of the AC Stark effect in atoms it is known that for a red-detuned beam the polarizability $\alpha$ is positive and for blue-detuned one it becomes negative. Similar conclusions can be drawn from a purely classical model of the atom [47, 48].
@@
The Newton equation of motion of an atom in these conditions takes the form
$m\ddot{\bm{r}}=({\bm{d}}\cdot{\bm{\nabla}}){\bm{E}}=\frac{1}{2}\,\alpha{\bm{\nabla}}(E^{2}),$ (24)
the right hand side of the equation being treated as averaged over the fast oscillations of the field in (2). The smoothed equations of motion can be given the following form in dimensionless coordinates $\xi_{x},\xi_{y},\zeta$, defined in (6):
@@
$\ddot{\bm{\zeta}}=\tilde{\gamma}\,\partial_{\zeta}|\psi(\bm{\xi},\zeta)|^{2},$ (25c)
@@
$\tilde{\gamma}=\frac{\alpha E_{0}^{2}}{4z_{R}^{2}\omega^{2}m}=\gamma\left(\frac{2}{kw_{0}}\right)^{2},$ (26b)
where $\omega=kc$.
@@
Depending on the sign of $\alpha$ the tubes described in the preceding section constitute either some kind of potential “valleys” (for $\alpha<0$) or repulsive-potential “hills” (for $\alpha>0$). Naturally, it is impossible to derive analytical solutions to the equations of motion in this kind of potentials, but trajectories of particles can be found numerically, for certain illustrative parameter values. In Fig. 7 it can be observed, from two different perspectives, that the trajectories of three negatively polarizable particles, having been inserted into the beam shown in Figs. 1, 2 and 5, follow the irradiance “holes”
Fig. 8 demonstrates the same phenomenon for five particles placed in the five-hole beam of Figs. 3, 4 and 6. For more complex beams, with a larger number of potential valleys, the issue becomes more challenging, owing to the very complicated arrangement of valleys and hills, which can result in jumping of particles between the tubes. The probability of such phenomenon to occur was established numerically for complex knotted vortex lines [43].
@@
As we know, if a particle with $\alpha>0$ is inserted into a beam of this type, regions of low irradiance should exert a repulsive effect on it. As a result, there ought to remain characteristic holes in the chaotic trajectory of such a particle, which cannot be penetrated. This situation is presented in Fig. 9 for $n=3$ and $n=5$ in the form of a projection of the calculated trajectories on the plane $\zeta=\mathrm{const}$. A three-dimensional drawing would be unreadable for obvious reasons.
The last figure of this section, i.e 10, illustrates the trajectories of two particles: one with $\alpha<0$ and the other with $\alpha>0$ placed simultaneously in a beam with $n=3$. As can be observed, the former is moving inside the hole created by the trajectory of the latter.
@@ IV Concluding remarks
In conclusion, it should be stressed that the choice of a suitable Gaussian beam prefactor gives the possibility of designing beams that exhibit low-intensity tubes. In Section II, the prefactor was chosen in the form of a polynomial corresponding to the superposition of two Gaussian beams: a fundamental one, and one that exhibits a vortex of the $n$th degree on the propagation axis. Due to the interference, this vortex gets splitted into $n$ vortices located symmetrically on the circle, resulting in the appearance of the mentioned black wormholes. They can constitute independent lines for guiding particles.
@@
The results of numerically performed calculations, presented in Section III, show that particles with negative polarizability, neglecting some transverse oscillations, do indeed move along trajectories determined by lines of vanishing wave intensity.
Trajectories calculated for positively polarizable particles show the opposite character: they avoid the mentioned areas.

### v2

@@ Abstract
The present paper discusses the application of certain special Gaussian beams that, thanks to some polynomial prefactors, have uniquely designed holes in the irradiance. Such holes, or rather tubes, can constitute potential valleys for negatively polarizable particles, providing the possibility of guiding several objects of that kind, each along its own trajectory. The mechanism of creating these holes by interference of Gaussian beams which exhibit orbital angular momentum is discussed, and then the trajectories of particles moving in such a wave are numerically calculated. As it turns out, these particles, performing transverse oscillations, follow the designed tunnels of low irradiance. On the contrary, for particles with positive polarizability these areas are inaccessible.
@@ I Introduction
The original optical tweezers operated due to gradient forces pulling particles of positive polarizability into areas of high wave intensity. In the similar way, the irradiance “holes” can serve as traps or guide lines for objects of negative polarizability such as atoms in the blue-detuned beams [16, 17, 18]. The identical effect is owed to the ponderomotive force acting on charged particles, such as electrons, originating from the inhomogeneous circularly polarized wave [19, 20].
As a recent experimental example of the use of so-called Dark Focus Tweezers one can recall the work [21] which demonstrates trapping of nanometer-sized silicon balls placed in an external medium to achieve a relative refractive index less than one. A hollow beam called a “bottle beam” [22] was obtained in this study by superposition of Gaussian and Laguerre-Gaussian beams. Numerous references to other theoretical and experimental findings can be found in the works cited above.
Dark Focus Tweezers or hollow beams have an advantage that particles are trapped or guided in the low radiation domain, and are therefore less likely to be damaged. This is particularly significant for biological objects.
@@
This type of beams has been known in the literature for about 30 years, although it seems that not in the context of guiding particles. To mention just a few results one can first recall the work [39], which, after a rather general introduction containing theoretical foundations of the so called arrays of vortices, provides numerical results concerning vortex-vortex or vortex-antivortex interactions. In turn the experimental work [40] demonstrates the generation of a two-hole beam thanks to the diffraction on the computer-synthesized grating and examins the stability of nearly located vortices. In [41] the authors focus on the numerical analysis of the motion and interaction of vortices inserted in a Gaussian beam, in a linear and nonlinear media.
This kind of structured light is expected to find potential application in communication. For this reason, the stability of hollow beams subjected to random phase distorsions has also been investigated [42].
@@ II Description of multi-hole Gaussian beams
In the above formulas (9) and (10) the complex coordinate $\xi=\xi_{x}+i\xi_{y}$ is introduced and should be distinguished from $|\bm{\xi}|=\sqrt{\xi_{x}^{2}+\xi_{y}^{2}}=|\xi|$. Result (10) is well known [52]. In particular the choice of $\tilde{\psi}(s)=\mathrm{const}$, leads to the fundamental GB, and $\tilde{\psi}(s)=\mathrm{const}\cdot s^{n}$ corresponds to the GB of vorticity $n$.
@@
$\left\{\begin{array}[]{l}|\bm{\xi}|=\displaystyle\frac{p_{0}}{\beta}\sqrt{1+\zeta^{2}},\\ \phi(\zeta)=\phi_{0}+\arctan\zeta,\end{array}\right.$ (11)
where $\beta$ is a certain constant playing the role of the scaling factor defined below in (12). Since all possible zeros follow synchronized, identical paths, they never merge, their number remains constant along the beam, and each develops its own nodal line as the value of $\zeta$ increases (naturally in the opposite direction, i.e. for $\zeta<0$ as well). This effect was observed in [39].
@@
The scaling role of the parameter $\beta$ has already been mentioned and is clear in (12). This form indicates that the interference of exactly two Gaussian beams is dealt with: that of order 00:
@@
Beam (12) bears the vortex topological charge $n$ and when encircling the axis $\zeta$ the value of the phase increases by $2\pi n$ which means that it assumes $n$ times the same values (if reduced to the interval $[0,2\pi[$). Thus, on a circle (for $\zeta=\mathrm{const}$) of radius $\sqrt{1+\zeta^{2}}/\beta$ (in units of $w_{0}$) a completely destructive interference with $\tilde{\psi}_{0}$ occurs exactly $n$ times. The $n$th degree vortex “spreads” into $n$ individual vortices (in the case of Bessel beams the same phenomenon was demostrated in [53]), uniformly distributed, as is obvious from the distribution of the $n$th complex roots of the unity. Consequently, there appear $n$ “holes” in the wave intensity in the perpendicular plane, as shown in Fig. 1 for $n=3$. In the subsequent diagrams (a)-(h) performed for increasing values of $\zeta$, the diffraction of the wave and the twisting of the whole pattern can be seen, in agreement with (11).
@@
Figures 3 and 4 demonstrate the same effects for $n=5$. The existence of five distinct roots of unity yields five “tubes” of vanishing irradiance that can eventually be exploited. A slightly reduced value for the parameter $\beta$ has been chosen in this case in order to avoid merging areas of low energy density.
@@ III Guiding of particles
As is well known, inhomogeneities in the intensity of the wave, and hence in the electric field, can be exploited to trap and guide neutral particles, for instance atoms or dielectric balls, which undergo polarization in external fields. This has become the basis for the operation of the so-called optical tweezers [12, 13, 14, 15], as mentioned in Introduction.
Let us denote with $\alpha$ the particle polarizability, which in general can depend on the driving frequency, and with $d$ the induced dipole moment. Then
@@
When speaking of atoms, from the theory of the AC Stark effect it is known that for a red-detuned beam the atomic polarizability $\alpha$ is positive and for blue-detuned one it becomes negative, which affects the motion of atoms in a fundamental way. Similar conclusions can be drawn from a purely classical model of the atom [55, 56]. The analogous efect in the case of dielectric nano-spheres is owed to the value of the refractive index higher or lower than that of the surrounding media.
@@
The Newton equation of motion of an atom (or other particle) in these conditions takes the form
$m\ddot{\bm{r}}=({\bm{d}}\cdot{\bm{\nabla}}){\bm{E}}=\frac{1}{2}\,\alpha{\bm{\nabla}}(E^{2})+\bm{F}_{\mathrm{scat}},$ (24)
the right hand side of the equation being treated as averaged over the fast oscillations of the field in (2). The first term represents the conservative gradient force and $\bm{F}_{\mathrm{scat}}$ stands for the scattering force to be dealt with below in some detail.
The irradiance in this kind of beams varies rapidly in directions orthogonal to the beam axis and relatively slowly alongside it, so the perpendicular components of $\bm{F}_{\mathrm{scat}}$ can be ignored compared to the gradient force and only parallel one needs to be accounted for, which is confirmed below. In order to estimate the magnitude of this force we will focus on the example of dielectric nano-spheres. The influence of this force on atoms can be reduced by using far-detuned beam, although it should be kept in mind that the gradient force is also weakened in this way.
As it is known [57, 58] for dielectric spheres one has
$\frac{F_{\mathrm{scat}\,z}}{F_{\mathrm{grad}\,z}}=\frac{n_{s}k^{4}\alpha}{3\pi n_{m}^{3}\varepsilon_{0}}\,\frac{I}{\partial_{z}I},$ (25)
where $n_{s}$ is the refractive index of the sphere, $n_{m}$ – the refractive index of the surrounding media, $\varepsilon_{0}$ – the vacuum permittivity and $I$ stands for the irradiance. In dimensionless coordinates $\xi_{x},\xi_{y},\zeta$ defined in (6) this ratio can be given the form
$\frac{F_{\mathrm{scat}\,\zeta}}{F_{\mathrm{grad}\,\zeta}}=\frac{n_{s}}{n_{m}^{3}}\,\zeta_{R}\kappa\,\frac{I}{\partial_{\zeta}I}$ (26)
with $\zeta_{R}=kz_{R}$ and $\kappa=\frac{k^{3}\alpha}{3\pi\varepsilon_{0}}$. Let us estimate these quantities for Rayleigh particles placed in a tightly focused light beam with the waist radius $w_{0}=0.05\,\mathrm{mm}$ and wavelength $\lambda=700\,\mathrm{nm}$ for which $\zeta_{R}\approx 10^{5}$. In order to find the magnitude of $\kappa$ one has to think of a concrete sample of particles. Take, for example, magnesium fluoride nano-balls ($n_{s}=1.38$) of diameter $D=10\,\mathrm{nm}$ placed in the glycerol ($n_{m}=1.47$). In these conditions one gets $\kappa\approx-1.1\cdot 10^{-5}$, where the well known formula for the polarizability [58] has been used:
$\alpha=\frac{1}{2}\,n_{m}^{2}\varepsilon_{0}D^{3}\,\frac{n_{s}^{2}/n_{m}^{2}-1}{n_{s}^{2}/n_{m}^{2}+2}.$ (27)
Consequently the whole factor in (26) equals
$\frac{n_{s}}{n_{m}^{3}}\,\zeta_{R}\kappa\approx-0.48,$ (28)
As regards the quotient $\frac{I}{\partial_{\zeta}I}$, it should be kept in mind that inside the “tube” of low irradiance, $I$ drops to zero (together with the scattering force). Naturally, the same refers to $\partial_{\zeta}I$ (since the vortex core constitutes the irradiation minimum) but the latter decline is weaker, as it is shown below.
For the beams dealt with in this work, defined in (12), (13) and (14) the needed ratio can be calculated to be
$\left|\frac{I}{\partial_{\zeta}I}\right|$ $=$ $\left|\frac{|\Psi|^{2}}{\partial_{\zeta}|\Psi|^{2}}\right|=\Bigg{|}-\frac{2(n+1)\zeta}{1+\zeta^{2}}+\frac{4|\xi|^{2}\zeta}{(1+\zeta^{2})^{2}}$ (29)
$-\sum_{l=0}^{n-1}\left(\frac{\partial_{\zeta}\xi_{l}}{\xi-\xi_{l}}+\frac{\partial_{\zeta}\xi_{l}^{*}}{\xi^{*}-\xi_{l}^{*}}\right)\Bigg{|}^{-1},$
where
$\xi_{l}=\frac{1}{\beta}(1+i\zeta)e^{\frac{2\pi il}{n}},\quad\mathrm{and}\quad\partial_{z}\xi_{l}=\frac{i}{\beta}\,e^{\frac{2\pi il}{n}}.$ (30)
For particles guided in the $i$th tube $|\xi-\xi_{i}|\ll 1/\beta$ and the $i$th term in the sum in(29) dominates over all other ones. Therefore
$\frac{F_{\mathrm{scat}\,\zeta}}{F_{\mathrm{grad}\,\zeta}}\approx\frac{n}{n_{s}^{3}}\,\zeta_{R}\kappa\,\left|\frac{\xi-\xi_{i}}{\partial_{\zeta}\xi_{i}}+\frac{\xi^{*}-\xi_{i}^{*}}{\partial_{\zeta}\xi_{i}^{*}}\right|\ll 1,$ (31)
and the scattering force should not have a relevant effect on the motion of guided particles.
This estimate has been carried out for tightly focused beams. They can be broadened at the price of increasing the scattering force (it grows quadratically with $w_{0}$ provided the local energy density is maintained). However, this has no significant effect on the trajectories of guided particles (there is a quantitative effect, but not a qualitative one), as tested numerically.
The scattering force has a little more significant effect for wider trajectories that depart from the core of vanishing irradiance but becomes truely essential for particles of positive polarizability which avoid areas of low irradiance. The motion of this type of particles is also shown in several figures below, therefore $F_{\mathrm{scat}\,\zeta}$ is taken into account in all numerical calculations.
Figure 7: The trajectories of three particles of negative polarizability placed in the “holes” of the beam in question for $n=3$, viewed from the side and from above. Straight lines represent zero irradiance tubes, which undergo diffraction. The values of beam parameters are the same as in Fig. 1, and $\gamma=6\cdot 10^{-2}$ and $\tilde{\gamma}=2\cdot 10^{-5}$.
Consequently the smoothed (with respect to time) equations of motion can be given the following form:
@@
$\ddot{\bm{\zeta}}=\tilde{\gamma}\left(\partial_{\zeta}|\psi(\bm{\xi},\zeta)|^{2}+\frac{n_{s}}{n_{m}^{3}}\,\zeta_{R}\kappa|\psi(\bm{\xi},\zeta)|^{2}\right),$ (32c)
@@
$\tilde{\gamma}=\frac{\alpha E_{0}^{2}}{4z_{R}^{2}\omega^{2}m}=\gamma\left(\frac{2}{kw_{0}}\right)^{2}=\frac{2}{\zeta_{R}}\,\gamma,$ (33b)
ande $\omega=kc$. Depending on the sign of $\alpha$ the tubes described in the preceding section constitute either some kind of potential “valleys” ($\gamma<0$) or repulsive-potential “hills” ($\gamma>0$). Naturally, it is impossible to derive analytical solutions to the equations of motion in this kind of potentials, but trajectories of particles can be found numerically, for certain illustrative $\gamma$-parameter value.
In Fig. 7 it can be observed, from two different perspectives, that the trajectories of three negatively polarizable particles, having been inserted into the beam shown in Figs. 1, 2 and 5, follow the irradiance “holes”. The value of $\gamma=0.06$ was chosen for visualization purposes. Note the expansion of the particle’s trajectory as it moves along the potential tube. It is related to the lowering of the potential barrier, which is associated with the beam’s diffraction and with some acceleration of the particle. This potential barrier decreases proportionally to $1/w(z)^{2}$ or to $1/(1+\zeta^{2})$. Over a distance of one Rayleigh length it then decreases twice. The potential valley becomes shallower, and the trajectory proportionally wider. From the practical point of view, such a distance seems sufficient.
Figure 8: Escape of a particle from the potential tube. The values of parameters are the same as in Fig. 7.
The lowering of the barrier height and width may be of importance for the eventual hopping of particles between different potential valleys. Especially in the quantum case, it affects the tunneling probability of guided elementary particles between various minima. Roughly speaking, if one neglects the possible acceleration of particles due to the gradient and scattering forces along the beam, the tunneling probability $p$ increases with $\zeta$ as $p^{(1+\zeta^{2})^{-1/2}}$.
Figure 8 shows the exemplary effect of the escape of one of the guided particles due to insufficient cooling and imprecise initial conditions which entail more chaotic motion. In [59] strong correlation between the chaoticness of classical trajectories of particles and the quantum tunneling probability in a trap constructed of a Bessel beam and constant magnetic field was established.
@@
Fig. 9 demonstrates the same phenomenon as Fig. 7 for five particles placed in the five-hole beam of Figs. 3, 4 and 6. For more complex beams, with a larger number of potential valleys, the issue becomes more challenging, owing to the very complicated arrangement of valleys and hills, which can result in jumping of particles between the tubes. The probability of such phenomenon to occur was established numerically for complex knotted vortex lines [50].
@@
As we know, if a particle with $\alpha>0$ is inserted into a beam of this type, regions of low irradiance should exert a repulsive effect on it. As a result, there ought to remain characteristic holes in the chaotic trajectory of such a particle, which cannot be penetrated. This situation is presented in Fig. 10 for $n=3$ and $n=5$ in the form of a projection of the calculated trajectories on the plane $\zeta=\mathrm{const}$. A three-dimensional drawing would be unreadable for obvious reasons. In line with the previous discussion, to weaken the effect of scattering force on the trajectory it was necessary to reduce the size of the spheres to $8\,\mathrm{nm}$.
The last figure of this section, i.e 11, illustrates the trajectories of two particles: one with $\alpha<0$ and the other with $\alpha>0$ placed simultaneously in a beam with $n=3$. As can be observed, the former is moving inside the hole created by the trajectory of the latter.
@@
This figure highlights the extremely chaotic nature of the positive-polarizable particle trajectories resulting from the eminently nonlinear form of the equations of motion in a very complex potential with many mninima, maxima and saddle points, versus fairly regular trajectories of negatively polarizable particles (the potential inside low irradiance tubes is quite smooth).
@@ IV Concluding remarks
In conclusion, it should be stressed that the choice of a suitable Gaussian beam prefactor gives the possibility of designing beams that exhibit low-intensity tubes. In Section II, the prefactor was chosen in the form of a polynomial corresponding to the superposition of two Gaussian beams: a fundamental one, and one that exhibits a vortex of the $n$th degree on the propagation axis. Due to the interference, this vortex gets splitted into $n$ vortices located symmetrically on the circle, resulting in the appearance of the mentioned black wormholes. They can constitute independent lines for guiding particles. The results of numerically performed calculations, presented in Section III, show that particles with negative polarizability, neglecting some transverse oscillations, do indeed move along trajectories determined by lines of vanishing wave intensity. The influence of the scattering force on the motion has been estimated in the case of dielectric nanospheres and it has not proved to exert a significant effect on them.
Trajectories calculated for positively polarizable particles show the opposite character: they are of very chaotic nature but avoid the mentioned areas. In this case, the scattering force should be accounted for, since its impact on trajectories is significant.